OpenStudy (sonyalee77):

Solve the system of equations by graphing. y = 1/2x + 3 y =4 x + 10

OpenStudy (johnweldon1993):

I hae graphing....but an algebraic approach to it...is see that y = two different things? well that must mean that those two things are equal as well... so we can do this 1/2x + 3 = 4x + 10 now just get all the like terms on one side 1/2x - 4x = 10 - 3 1/2x - 8/2x = 7 -7/2x = 7 now multiply both sides by -2 7x = -14 divide by 7 x = -2 :)

OpenStudy (anonymous):

y = 2

OpenStudy (sonyalee77):

but how do u graph it

OpenStudy (johnweldon1993):

@Soumyadeep ..if it were 2..the top equation would = 4 and the bottom would = 18

OpenStudy (johnweldon1993):

and to graph it...it would look something like

OpenStudy (johnweldon1993):



OpenStudy (anonymous):

Sorry boss! didnt check the other one!

OpenStudy (johnweldon1993):

Actually I had 2 at first too lol just had to recheck myself and say the mistake